OpenStudy (anonymous):

The value of tan^-1[tan 5pie/6

OpenStudy (anonymous):

tan^-1[tan 5pi/6] ?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

this might come in handy \[\huge \tan^{-1} [\tan x] = x\]

OpenStudy (lgbasallote):

that'snot always the case but usually it is

OpenStudy (anonymous):

the answer should be -pie/6

OpenStudy (anonymous):

@satellite73

terenzreignz (terenzreignz):

\[\tan^{-1}[\tan x] = x + k\pi \] for as long as k is an integer?

OpenStudy (lgbasallote):

that's simplified form i believe

OpenStudy (anonymous):

@yahoo! what are your options

OpenStudy (shubhamsrg):

tan^-1 (tan x) = x where x lies in the range of (-pi/2 , pi/2) =>5pi/6 is out of that range we need to re arrange note that tan(5p/6) = tan(pi - pi/6) = -tan(pi/6) now -tan(pi/6) lies in that range so you see why thats the ans..

OpenStudy (anonymous):